### v1

@@ 3.2.1 F1-score
Accuracy: Several subsets of our benchmark – Presence, Dimension, and Functional Performance questions – ask the model to provide a yes/no answer. We score these using accuracy (ACC).
3.2.2 BLEU
@@
The BLEU (Bilingual Evaluation Understudy) score was developed by Papineni et al. [31] for the automatic scoring of machine translations relative to human, reference translations. It has been employed in recent benchmarks [22, 32] and was used by [22] to evaluate the semantic similarity between an explanation provided by a generative model and an explanation written by a human. To compute BLEU, predicted and reference sentences are first broken up into n-grams, which are segments of n (a user-specified number) words. n-gram matches between the predicted and reference sentences are then found; once a predicted n-gram is matched to a reference n-gram, the n-gram is removed from the pool of reference n-grams in a process called ‘‘clipping."The number of matching n-grams is then divided by the number of n-grams in the predicted sentence, producing a modified precision score, $p_{n}$. The authors suggest to compute BLEU-4: $p_{n}$ for n = 1 through n = 4, taking the geometric mean of the four $p_{n}$ (since $p_{n}$ decays exponentially with increasing n) [31]. However, since we only use one reference (one explanation), BLEU-4 scores were always near zero. As such, we report BLEU-2, which has non-zero scores but still preserves some of the information about adjacency of words. BLEU-2 (with max n-gram 2) can be computed as: $\log(\text{BLEU})=\min\left(1-\frac{r}{c},0\right)+\sum_{n=1}^{2}\frac{1}{2}\log(p_{n})$
@@
The min() term is a ‘‘brevity penalty"and serves to penalize predictions (of length c) that are shorter than the ground truth (length r). For the explanation portions of the Rule Compliance questions, we report BLEU-2 to quantify the similarity between the model’s generated explanation and the human-generated explanation.
3.2.3 ROUGE
@@
The ROUGE (Recall-Oriented Understudy for Gisting Evaluation) metric was developed by Lin et al. [33] to measure the quality of a computer-generated text summary relative to a human, reference summary. The metric has been utilized in recent benchmarks [22, 14] to assess generative models’ abilities to craft explanations and summarize text. ROGUE-L, which uses longest common sub-sequence (LCS), is useful for characterizing similarity in sentence-level word order. The LCS refers to the longest sequence of words that appear in the same order but not necessarily consecutively in the two sentences. ROGUE-L is computed as the F1 score of the LCS: $R_{LCS}=\frac{LCS(R,P)}{m}$, $P_{LCS}=\frac{LCS(R,P)}{n}$, and $ROUGE_{L}=\frac{2P_{LCS}\times R_{LCS}}{P_{LCS}+R_{LCS}}$,
@@
Таблица 2: This table presents a detailed comparison of various MLLM models on our dataset, revealing GPT4’s superior performance of the models tested. The results underscore that despite the advancement of these models in understanding and applying complex technical specifications, there is significant room for improvement for AI models to understand engineering requirements.
Evaluation Metric Baseline Model
Section Subset (Metric) Naive GPT4-AllRules GPT4-RAG GPT4-NoRules LLaVA-RAG LLaVA-NoRules
Extraction Retrieval (F1 BoW) 0.082 0.750 0.181 N/A 0.112 N/A
Compilation (F1 rules) 0.137 0.298 0.362 N/A 0.281 N/A
Comprehension Definition (F1 BoC) 0.358 0.470 N/A 0.420 N/A 0.393
Presence (ACC) 0.5 0.629 0.532 N/A 0.484 N/A
Compliance Dimension (ACC/Bleu/Rouge) 0.5/-/- 0.533/0.118/0.296 0.300/0.091/0.235 N/A 0.408/0.097/0.241 N/A
Functional Performance (ACC/Bleu/Rouge) 0.5/-/- 0.563/0.167/0.342 0.563/0.121/0.306 N/A 0.536/0.163/0.321 N/A
@@ 4 MODEL EVALUATION
We evaluate simple baselines and recent state-of-the-art MLLMs on our DesignQA benchmark to understand the state of current AI models in understanding engineering requirement documentation. The goal is to identify gaps in current AI’s capabilities and encourage other researchers to build and train better AI models and frameworks for answering the questions we collected. Moreover, evaluating different MLLM models provides some insight into the relative difficulty of the questions, and the failure modes could serve as inspiration for better approaches to the benchmark.
@@ 4.1 Baselines and Models
MLLM Models
@@
We consider two recent MLLMs in our evaluation: the closed-source model from OpenAI gpt-4-1106-vision-preview (GPT4) and the open-source model llava-1.5-13b (LLaVA) [1, 34]. GPT4 is chosen for evaluation because of its high performance on existing benchmarks [14, 9] while LLaVA is selected because of its promise (and the promise of its derivatives) as an open-source MLLM [35]. The extracted text from the FSAE rule document PDF is roughly 70,091 tokens in length. While GPT4 has a 128,000 token context window and can ingest the whole text in the prompt, LLaVA only has a 4,096 token context window and cannot. Thus, LLaVA requires the use of a Retrieval Augmented Generation (RAG) system to retrieve the appropriate context relevant to the question from the FSAE rule document. RAG is a natural language processing technique that enhances text generation by incorporating external knowledge, dynamically retrieving relevant information from a database or corpus to inform and improve the content being produced. It is especially helpful for models with small context windows, as it selectively pulls relevant information from an extensive document to assist in generating more informed and contextually accurate text outputs. While we need RAG to test the open-source LLAVA model, our goal is not to study different RAG techniques. To this end, we implement a very simple RAG system using LlamaIndex and using OpenAI’s text-embedding-3-large to embed the information in the FSAE rule document [36]. We index the text from the rule document into 250-token chunks, with a 50-token overlap. From the embeddings, the cosine similarity between the question and each of the embedded chunks is then computed, and the top-15 (top-12 for Compliance QAs) most relevant pages are then fed into the prompt as context, before posing the question in the benchmark.
While there are few images in the FSAE rule document, they were not considered for our evaluation, as they require further processing, and LLaVA was not trained on multiple image inputs. The few tables in the FSAE rule document will not receive any special treatment and will be fed into the models just as simple text from the PDF text-extracting script, together with the rest of the text on the page.
For all subsets of the benchmark, except the Definition questions, we evaluate the performance of GPT4 given the full rule document via its context window (GPT4-AllRules), GPT4 given the rule document via the LlamaIndex simple RAG (GPT4-RAG), and LLaVA given the rule document via the LlamaIndex simple RAG (LLaVA-RAG). While the Retrieval and Compilation questions in the dataset could be answered with text-only models, we pose the questions in this segment with a null image to the MLLMs. For the Definition questions, RAG always returns the same portion of the rule document since the questions across each QA are the same (i.e., ‘‘tell me the name of the highlighted component,"even though the images vary across QAs). Because of this, we do not test GPT4-RAG and LLaVA-RAG on the Definition questions, but we instead test GPT4 and LLaVA without any input of the rules (GPT4-NoRules, LLaVA-NoRules) in addition to GPT4-AllRules.
@@ 4.2 Results and Analysis
For all subsets of the benchmark, models performed better than the naive baseline. While no single model performed the best across all subsets of the benchmark, GPT4-AllRules was the dominant model in all subsets except the Compilation questions, where GPT4-RAG was the best performer. This finding reflects that the simple LlamaIndex RAG model used was not very effective at providing relevant rule information to the model, while inputting the full rule text into the context window allowed GPT4-AllRules to access necessary information. For the Extraction and Comprehension segments of the benchmark, GPT4 (-RAG or -NoRules) performed better than the corresponding LLaVA model. The reverse was true for the Compliance Segment: LLaVA-RAG was equivalent to or outperformed GPT4-RAG.
@@
Of the models tried, GPT4-AllRules does the best job at retrieving the requested rule verbatim (0.750 average BoW score). When the model’s answer diverged from the ground truth answer, we noticed that it was often because the model was reporting a nearby rule rather than the rule requested (e.g. V.3.2.5 instead of V.3.2.4). Sometimes GPT4-AllRule’s answer was different from the ground truth because it included all child rules in addition to the requested parent rule (e.g when asked for V.1, V.1 was reported along with V.1.1 and V.1.2). While not technically wrong, this result was not handled differently by our evaluation metric and may be a result we would want to handle specifically in the future.
Providing the rules to GPT4 via RAG instead of via context resulted in a much lower average BoW score (0.181 for GPT4-RAG). For a number of questions (like the GPT4-RAG Retrieval example shown in Figure 3) the model replied that it could not produce the text for the rule because it was not included in the rule text given to it, indicating that the simple LlamaIndex RAG failed to provide the relevant portion of the rule document to the model. When GPT4-RAG did provide an answer to the question, it had similar responses to that of GPT4-AllRules. LLaVA-RAG received the same portion of the rule document as GPT4-RAG, but its average BoW score was even lower (0.112). Unlike GPT4-RAG, LLaVA-RAG would hallucinate rules rather than indicate that the requested rule was not contained within the portion of the document it received via RAG. Furthermore, instead of returning rules verbatim, LLaVA-RAG would frequently offer an interpretation of the requested rule (as in the LLaVA-RAG Retrieval example in Figure 3).
@@
GPT4-RAG performs the best at the Compilation questions (0.362 F1 rules). We noticed that GPT4-RAG returns many fewer rules per question on average (9.57 rules) than GPT4-AllRules (17.53 rules), as seen in Figure 3. This could explain the improved performance of GPT4-RAG relative to GPT4-AllRules: when given the full 140-page rule document, the model tends to over-predict relevant rules, resulting in low precision and reducing the average F1 score. When given the top-15 most relevant pages via RAG, the model has fewer rules that it can draw from and therefore does not include as many false positives in its responses. LLaVA-RAG has the lowest score of the models tested (0.281). For five out of the 30 questions, its predicted rule list has a score of zero (no overlap with the ground truth rule list). In contrast, GPT4-RAG has a score of zero for two out of the 30 questions.
Таблица 3: Performance analysis of MLLM models on definition QA, highlighting the impact of component mention types within rule documents. GPT4-AllRules leads in direct definition mentions, while all models show varied effectiveness across multi-mention and no-mention scenarios, underlining the challenges in extracting definitions without explicit mentions.
F1 BoC GPT4 -AllRules GPT4 -NoRules LLaVA -NoRules
Definition
0.74 0.56 0.50
Multi-mention
0.42 0.39 0.37
No-mention
0.35 0.35 0.37
@@
As explained in Section 4.1, -NoRules versions of GPT4 and LLaVA were used for this subset of the benchmark rather than -RAG versions, since the identical question text across QAs resulted in RAG always incorrectly returning the same section of the rule text. Of the models tested, GPT4-AllRules has the highest score (0.470 average F1 BoC). However, GPT4-NoRules performs better than we expected (0.420 average F1 BoC), indicating that identification of some components in an FSAE vehicle can be performed with knowledge from the web. As explained in Section 3.1.3, we tracked how the component-in-question in these QAs was mentioned in the rule document (definition-component, multi-mention component, or no-mention component). As to be expected, the drop in F1 BoC score between GPT4-AllRules and GPT4-NoRules is due mostly to a drop in score for the definition-component QAs (Table 3), which without the definitions spelled-out in the rule document, become harder to answer.
@@
GPT4-AllRules performed the best (0.629 accuracy) on the Presence questions of the models tested. GPT4-RAG has a lower average accuracy (0.532), likely because the model has access to limited (15 or fewer) rule document chunks that reference the component in question. Like the Dimension questions, LLaVA-RAG performs marginally worse than GPT4-RAG.

### v2

@@ 3.2.2 Accuracy
Several subsets of our benchmark – Presence, Dimension, and Functional Performance questions – ask the model to provide a yes/no answer. We score these using accuracy (ACC).
3.2.3 Explanation Metrics
For the Rule Compliance questions, models return explanations as well as yes/no answers. Automatically scoring model-generated explanation text for semantic similarity to a human-written reference text is a complex task, with no perfect solution. A recently developed benchmark testing large language models on multiple-choice science questions, ScienceQA [25], used three different automated metrics – BLEU [37], ROUGE [38], and Similarity [39] – to assess similarity between LLM-generated and human-written explanations. We follow Lu et al.’s example and use these three metrics to score the MLLM-generated explanations relative to the human-collected ones in our benchmark. Each of these metrics is described further below.
@@ 3.2.3 Explanation Metrics
BLEU: The BLEU (Bilingual Evaluation Understudy) score was developed by Papineni et al. [37] for the automatic scoring of machine translations relative to human, reference translations. To compute BLEU, predicted and reference sentences are first broken up into n-grams, which are segments of n (a user-specified number) words. n-gram matches between the predicted and reference sentences are then found; once a predicted n-gram is matched to a reference n-gram, the n-gram is removed from the pool of reference n-grams in a process called “clipping.” The number of matching n-grams is then divided by the number of n-grams in the predicted sentence, producing a modified precision score, $p_{n}$. The authors suggest to compute BLEU-4: $p_{n}$ for n = 1 through n = 4, taking the geometric mean of the four $p_{n}$ (since $p_{n}$ decays exponentially with increasing n) [37]. However, since we only use one reference (one explanation), BLEU-4 scores were always near zero. As such, we report BLEU-2, which has non-zero scores but still preserves some of the information about adjacency of words. BLEU-2 (with max n-gram 2) can be computed as: $\log(\text{BLEU})=\min\left(1-\frac{r}{c},0\right)+\sum_{n=1}^{2}\frac{1}{2}\log(p_{n})$
@@
The min() term is a “brevity penalty” and serves to penalize predictions (of length c) that are shorter than the ground truth (length r). For the explanation portions of the Rule Compliance questions, we report BLEU-2 to quantify the similarity between the model’s generated explanation and the human-generated explanation.
@@
ROUGE: The ROUGE (Recall-Oriented Understudy for Gisting Evaluation) metric was developed by Lin et al. [38] to measure the quality of a computer-generated text summary relative to a human, reference summary. ROUGE-L, which uses longest common sub-sequence (LCS), is useful for characterizing similarity in sentence-level word order. The LCS refers to the longest sequence of words that appear in the same order but not necessarily consecutively in the two sentences. ROUGE-L is computed as the F1 score of the LCS: $R_{LCS}=\frac{LCS(R,P)}{m}$, $P_{LCS}=\frac{LCS(R,P)}{n}$, and $ROUGE_{L}=\frac{2P_{LCS}\times R_{LCS}}{P_{LCS}+R_{LCS}}$,
@@
Similarity: The Similarity metric refers to Sentence-BERT-based sentence embedding cosine similarity [39, 40]. Both the model-generated explanation and the human-written explanation are embedded using the sentence-transformers/all-MiniLM-L6-v2 sentence transformer, resulting in two vectors that encode the meaning of the explanations. Cosine similarity between the two vectors is then computed to identify how much overlap there is in meaning between the two explanations.
@@ 4 MODEL EVALUATION
We evaluate simple baselines and state-of-the-art MLLMs, at the time of writing, on our DesignQA benchmark to understand current AI models’ capabilities in understanding engineering requirement documentation. The goal of these evaluations is not to be exhaustive: there are many other models that should be tested, time and cost permitting. Given the rapid pace of MLLM development, we ask for the research community’s assistance in benchmarking new state-of-the-art models as they are developed. The goal of the evaluations we conducted is rather to illustrate the gaps in current AI’s capabilities and encourage other researchers to build and train better AI models and frameworks that target these gaps. Moreover, evaluating different MLLM models provides some insight into the relative difficulty of the questions, and the failure modes could serve as inspiration for future approaches to benchmarking design-requirement-type questions.
@@ 4.1.1 Naive Baselines
4.1.2 MLLM Models
We consider five recently developed MLLMs in our evaluation: four closed-source models – OpenAI’s gpt-4o (GPT-4o) [16], OpenAI’s gpt-4-1106-vision-preview (GPT-4) [1], Google AI’s models/gemini-1.0-pro-vision (Gemini-1.0)  [^4] https://ai.google.dev/gemini-api/docs/models/gemini?authuser=1, and Anthropic’s claude-3-opus-20240229 (Claude-Opus)  [^5] https://www.anthropic.com/news/claude-3-family – and one open-source model: llava-1.5-13b (LLaVA-1.5) [17]. Note that all future references to GPT-4o, GPT-4, Gemini-1.0, Claude-Opus, or LLaVA-1.5 refer to these specific model versions unless otherwise stated. GPT-4, Gemini-1.0, and Claude-Opus were originally selected for their strong performances as closed-source models on existing multimodal benchmarks [14, 9]. For example, Gemini-1.0 Ultra, Claude-Opus, and GPT-4V(ision) were the top three models, as of April 2024, on the MMMU benchmark, which consists of 11.5k multimodal questions from college curriculum content [24]. GPT-4o was released towards the end of the time of this work. It holds the 1st place spot on the MMMU leaderboard [24], as of August 2024. It was added to our closed-source model evaluations both because of this notable performance and because of potential for comparison with GPT-4. In this way, DesignQA can be used to continue to measure progress of models as they develop over time. LLaVA-1.5 is selected because of its promise (and the promise of its derivatives) as an open-source MLLM [41]. While the Retrieval and Compilation questions in the benchmark could be answered with text-only models, we pose the questions in this segment with a null image to the MLLMs.
The extracted text from the FSAE rule document PDF is roughly 70,091 tokens in length. As GPT-4o and GPT-4 have 128,000 token context windows and Claude-Opus has a 200,000 token context window, these models can ingest the whole rule text in the prompt. However, LLaVA-1.5 and Gemini-1.0, which only have 4,096 token and 12,288 token context windows respectively, cannot take the whole rule text as context input. Thus, these two models require a Retrieval Augmented Generation (RAG) system to retrieve the appropriate context relevant to the question from the FSAE rule document [42]. As our research goal was not to evaluate different RAG techniques, we implement a very simple RAG using the LlamaIndex framework, described in more detail in Section 4.1.3. In the cases of GPT-4o, GPT-4, and Claude-Opus, ingesting the whole rule document via the context window proved to be costly, ten times or more the cost of providing the rule document to the model via the simple RAG. Of these models, Claude-Opus had the most expensive cost per input token at the time of writing, so we only tested Claude-Opus with a RAG system. For GPT-4o and GPT-4, we tested the models by providing them the full rule document via their context windows and also by providing them the rule document via RAG, so as to enable better comparison with the other models that received the rules via RAG.
In summary, we tested the following model combinations: GPT-4o-AllRules (given the full rule document via its context window), GPT-4-AllRules (given the full rule document via its context window), GPT-4o-RAG, GPT-4-RAG, LLaVA-1.5-RAG, Gemini-1.0-RAG, and Claude-Opus-RAG. For all models, we did not specify a system prompt template, so the default system prompt was used. We also used the default parameters (top_p and temperature) for each model, as set by LlamaIndex, the framework from which we called the MLLMs. More details can be found in the SI.
4.1.3 RAG
@@
RAG is a natural language processing technique that enhances text generation by incorporating external knowledge, dynamically retrieving relevant information from a database or corpus to inform and improve the content being produced [42]. It is especially helpful for models with small context windows, as it selectively pulls relevant information from an extensive document to assist in generating more informed and contextually accurate text outputs. As mentioned before, our goal in this work is not to study different RAG techniques. As such, we implemented a very simple RAG system using LlamaIndex and using OpenAI’s text-embedding-3-large to embed the information in the FSAE rule document [43]. We index the text from the rule document into 250-token chunks, with a 50-token overlap. From the embeddings, the cosine similarity between the question and each of the embedded chunks is then computed. The top-15 (top-12 for Compliance QAs) most relevant pages are then fed into the prompt as context, before posing the question in the benchmark. If including the top-15 most relevant pages exceeded the model’s context window size, the top-k value was lowered. While there are a few images in the FSAE rule document, they were not considered for our evaluation or included in our RAG, as they require further processing, and LLaVA-1.5 was not trained on multiple image inputs. The few tables in the FSAE rule document will not receive any special treatment and will be fed into the models just as simple text from the PDF text-extracting script, together with the rest of the text on the page.
For just the Definition question subset of the benchmark, the *-RAG models were not given any RAG-generated context. This is because the questions in the Definition subset are the same, i.e., “tell me the name of the highlighted component,” even though the images vary across QAs. As such, RAG always returned the same portion of the rule document for each question. Since this identical context is unhelpful, we did not provide it to the models for the Definition questions. As such, all *-RAG models tested did not receive document context just for the Definition subset.
@@ 4.2 Results and Analysis
Of the models tested, GPT-4o-AllRules performs the best across almost all metrics in the benchmark. GPT-4o-AllRules and GPT-4-AllRules, in all but one subset case, perform better than their corresponding -RAG models, demonstrating that the simple LlamaIndex RAG implementation was generally ineffective at providing relevant rule information to the models. In contrast, inputting the full rule text into the context window allowed GPT-4o-AllRules and GPT-4-AllRules to access necessary information. Of the models tested, GPT-4o-AllRules is not the best performer on the Rule Compliance explanations: Gemini-1.0-RAG receives higher scores (or ties) on the BLEU and ROUGE metrics, likely because of the model’s succinct explanations, and Claude-Opus-RAG ties GPT-4o-AllRules on the Functional Performance explanation Similarity score. Of the *-RAG models tested, GPT-4o-RAG tends to be the best performer, with the exception of the Functional Performance questions, where Claude-Opus-RAG has the highest Functional Performance accuracy score of the *-RAG models tested. This finding perhaps indicates that Claude-Opus has better engineering-design technical reasoning skills than the other models evaluated.
Despite GPT-4o-AllRules’s notable performance of the models tested, it does not attain perfect scores (1.0 for each subset of the benchmark). Even though it receives the full rule document, GPT-4o-AllRules cannot reliably extract specific rules from the text, a trivial yet critical task for engineers. MLLMs’ scores on the Compilation, Definition, and Presence subsets of the benchmark also remain relatively low when compared with the Naive baseline, possibly because these subsets’ questions don’t point the models to specific sections of the rule document, resulting in harder search problems. There remains significant work that must be done to develop MLLMs equipped with engineering and reasoning knowledge that can solve these problems characteristic of real-world engineering design challenges. We must focus efforts on developing MLLM methods that excel in engineering contexts, enhancing precision and reliability in real-world applications.
Table 2: This table presents a detailed comparison of various MLLM models’ scores on our benchmark. Of the models tested, GPT-4o-AllRules is the best performer across nearly all metrics. The Compilation, Definition, and Presence questions have the smallest difference between the best performing model’s score and the Naive baseline’s score. For all metrics, a perfect score would be 1.0 and
larger values are better, as indicated by the up arrows. Subset (Metric) Baseline Model Naive GPT-4o-AllRules GPT-4-AllRules GPT-4o-RAG GPT-4-RAG LLaVA-1.5-RAG Gemini-1.0-RAG Claude-Opus-RAG RULE EXTRACTION Retrieval (F1 BoW $\uparrow$) 0.082 0.885 0.750 0.186 0.181 0.112 0* 0.173 Compilation (F1 rules $\uparrow$) 0.137 0.424 0.298 0.376 0.362 0.281 0.283 0.288 RULE COMPREHENSION Definition (F1 BoC $\uparrow$) 0.358 0.540 0.470 0.525 0.420 0.393 0.488 0.423 Presence (ACC $\uparrow$) 0.5 0.726 0.629 0.710 0.532 0.484 0.548 0.5 RULE COMPLIANCE Dimension (ACC $\uparrow$) (BLEU/ROUGE/Similarity $\uparrow$) 0.5 - 0.825 0.175/0.344/0.777 0.533 0.118/0.296/0.728 0.675 0.111/0.255/0.642 0.300 0.091/0.235/0.592 0.408 0.097/0.241/0.578 0.525 0.176/0.344/0.644 0.508 0.137/0.295/0.698 Functional Performance (ACC $\uparrow$) (BLEU/ROUGE/Similarity $\uparrow$) 0.5 - 0.938 0.230/0.408/0.745 0.563 0.167/0.342/0.697 0.750 0.181/0.367/0.736 0.563 0.121/0.306/0.698 0.536 0.163/0.321/0.650 0.438 0.266/0.444/0.725 0.875 0.172/0.354/0.745 *Gemini-1.0-RAG received a score of 0 on the Retrieval question because of a “RECITATION” error that caused it to stop response generation.
@@
Of the models evaluated, GPT-4o-AllRules and GPT-4-AllRules significantly outperform the other models at retrieving the requested rule verbatim (0.885 and 0.750 average F1 BoW score, respectively). When an *-AllRules model’s answer diverged from the ground truth answer, we noticed that it was often because the model was reporting a nearby rule rather than the rule requested (e.g. V.3.2.5 instead of V.3.2.4) or because the model was reporting a similarly numbered rule (e.g., F.3.2.1) instead of the requested rule (e.g., V.3.2.1). Sometimes the *-AllRules models’ answers were different from the ground truth because they included all child rules in addition to the requested parent rule (e.g., when asked for V.1, V.1 was reported along with V.1.1 and V.1.2). While not technically wrong, this result was not handled differently by our evaluation metric and may be a result we would want to handle specifically in the future.
Providing the rules to GPT-4o or GPT-4 via RAG instead of via context resulted in a much lower average F1 BoW score (0.186 and 0.181, respectively). For a number of questions (like the GPT-4-RAG Retrieval example shown in Figure 3), GPT-4o-RAG and GPT-4-RAG replied that they could not produce the text for the rule because it was not included in the rule text given to it, indicating that the simple LlamaIndex RAG failed to provide the relevant portion of the rule document to the models. When GPT-4o-RAG and GPT-4-RAG did provide an answer to the question, they had similar responses to that of their respective *-AllRules models. Claude-Opus-RAG had a similar average F1 BoW score (0.173) to that of GPT-4o-RAG and GPT-4-RAG, and Claude-Opus-RAG also noted when the rule was not present in the provided RAG context.
LLaVA-1.5-RAG received the same portion of the rule document as the other *-RAG models, but its average F1 BoW score was even lower (0.112). Unlike GPT-4o-RAG, GPT-4-RAG, and Claude-Opus-RAG, LLaVA-1.5-RAG would hallucinate rules rather than indicate that the requested rule was not contained within the portion of the document it received via RAG. Furthermore, instead of returning rules verbatim, LLaVA-1.5-RAG would frequently offer an interpretation of the requested rule (as in the LLaVA-1.5-RAG Retrieval example in Figure 3).
Gemini-1.0-RAG received a score of zero on the Retrieval subset of the benchmark because the model encountered a “RECITATION” error for each question. The model has been programmed to stop content generation if the generated content repeats parts of the input data. While in some cases this behavior is useful, it is not useful for the purpose of this subset of the benchmark where we request verbatim repetition. We can imagine many other similar queries – where direct quotation is desired by a user – for which this repetition blocking would be problematic. Since this is a closed-source model, there is no way for us to alter this behavior.
@@
GPT-4o-AllRules performs the best (0.424 F1 rules) on the Compilation questions of the models tested. However, the Compilation questions are one of the benchmark categories (along with the Definition and Presence questions) which has a relatively small difference between the best performing model score (GPT-4o-AllRules) and the naive random baseline score. What inhibits GPT-4o-AllRules – the best performer – from having a higher score? We noticed that the model repeatedly fails to include rules in its answer that contain the search term verbatim. For example, in the first Compilation question, which asks the model for rules that pertain to “aerodynamic or aerodynamics,” the model neglects to include T.7.2.1, T.7.2.2, and T.7.5, which all include “aerodynamic” verbatim in their rule text. Improving MLLMs’ basic text extraction skills is essential to their performing better on design-requirement-type tasks. We also observed that GPT-4o-AllRules frequently neglects to report rules that are mentioned by rules that include the search term. For example, in the aerodynamic question, GPT-4o-AllRules doesn’t mention GR.6.4.1, V.1.1, and V.1.4.1, all of which are referenced by other rules that contain the word “aerodynamic.” While perhaps our prompt should explicitly state that rules referred to by rules that contain the search term are relevant, MLLMs must have the capability to follow chains of rules across disconnected portions of a document.
In terms of the *-RAG models, GPT-4o-RAG and GPT-4-RAG performed similarly on the Compilation questions, while the other *-RAG models didn’t perform as well. Claude-Opus-RAG’s lower score (compared with GPT-4o-RAG’s) could be attributed to its failure to follow the provided instructions. The prompt specifically asked for a list rules separated by commas, using no other words in the response. Claude-Opus-RAG includes other words in each of its answers, starting each answer with “The rules relevant…” and occasionally separating rules by spaces instead of commas. Since automated evaluations were used, this lack of instruction following resulted in lower scores. Gemini-1.0-RAG and LLaVA-1.5-RAG also had lower scores (0.283 and 0.281, respectively) than GPT-4o-RAG. The models’ answers to 6/30 (Gemini-1.0-RAG) and 5/30 (LLaVA-1.5-RAG) of the questions had a score of zero (no overlap with the ground truth rule list). In contrast, GPT-4o-RAG had a score of zero for just one out of the 30 questions.
Table 3: Results from the GuaranteedRAG experiment. We compare LLaVA-1.5’s performance on the benchmark when given FSAE rule document sections via smiple LlamaIndex RAG and when given GuranteedRAG.
Subset (Metric) Model
LLaVA-1.5-RAG LLaVA-1.5-GuaranteedRAG
RULE EXTRACTION
Retrieval (F1 BoW $\uparrow$) 0.112 0.699 $\pm$ 0.00590
Compilation (F1 rules $\uparrow$) 0.281 0.357 $\pm$ 0.0205
RULE COMPREHENSION
Definition (F1 BoC $\uparrow$) 0.393 0.471 $\pm$ 0.0255
Presence (ACC $\uparrow$) 0.484 0.543 $\pm$ 0.0353
RULE COMPLIANCE
Dimension (ACC $\uparrow$) (BLEU/ROUGE/Similarity $\uparrow$) 0.408 0.097/0.241/0.578 0.482 $\pm$ 0.0473 0.121 $\pm$ 0.008/0.266 $\pm$ 0.0105/0.652 $\pm$ 0.0158
Functional Performance (ACC $\uparrow$) (BLEU/ROUGE/Similarity $\uparrow$) 0.536 0.163/0.321/0.650 0.375 $\pm$ 0.159 0.152 $\pm$ 0.0165/0.316 $\pm$ 0.0269/0.650 $\pm$ 0.0154
@@
Of the models tested, GPT-4o-AllRules performs the best on the Definition subset of the benchmark. The Definition questions had the smallest difference between the best-performing model score and the naive random baseline score of any subset of the benchmark, indicating that they are challenging for MLLMs to answer. Contributing to GPT-4o-AllRules’s relatively low score on the Definition questions were visual component recognition errors. For example, when shown independent images where the steering wheel, steering column, and steering rack were each highlighted in pink, GPT-4o-AllRules never mentioned something related to “steering” as the highlighted component in-question. Instead, it answered “impact attenuator” twice and “front hoop” once. These components are very different parts of the vehicle from the steering system. As such, visual component recognition capabilities, especially when it comes to technical components, needs to be improved for MLLMs. Without good component understanding, MLLMs will struggle to answer Rule Compliance questions that pertain to those components.
@@
GPT-4o-AllRules performed the best (0.726 accuracy) on the Presence questions of the models tested. The *-RAG models all had lower average accuracies likely because these models had access to limited (15 or fewer) rule document pages that reference the component in question. Notably, Claude-Opus-RAG and LLaVA-1.5-RAG perform equivalent to and worse than the naive baseline, respectively. For 28/30 of the questions, Claude-Opus-RAG answered “no” to the posed question, perhaps reflecting a cautious stance. Since only half of the questions had a ground truth answer of no, consistently answering no amounts to the same as randomly guessing (the naive baseline). LLaVA-1.5-RAG had no noticeable trends in yes/no prediction, but performed worse than random choice.